# Provably Safe Reinforcement Learning using Entropy Regularizer

### v1

@@ 1 Introduction
To formalize this setup, we employ a CMDP formulation [3]. We assume that the target set is terminal, meaning that the process terminates whenever the MDP reaches a target state, while the unsafe and living sets are non-terminal. In contrast, our earlier work [15] assumed that the unsafe set was also terminal in order to reduce computational complexity.
@@ 2 Background and Problem Formulation
For $t\leq\tau_{E}$, we define a sequence $\{c_{t}\}$ with $c_{t}:=c(X_{t},A_{t})$. Similarly for $t\leq\tau$, we define another process $\{\kappa_{t}\}$ with $\kappa_{t}:=\kappa(X_{t},A_{t})$.
Definition 1.
@@
If the safe baseline policy is enforced only on the proxy states while an arbitrary (possibly exploratory) policy is applied elsewhere, the overall safety of the system remains guaranteed [17]. Hence, the proxy set enables enhanced exploration by allowing less conservative actions without compromising safety. This work is, however, not restricted to the prior knowledge of the proxy set. In scenarios where the proxy set is unknown or cannot be identified apriori, we adopt a conservative assumption by treating all states as proxy states.
@@
The computed policy ${\pi}_{k}(a|x)$ is kept fixed till the end of the episode $k$. Optimization (7)-(8) is based on the optimization in the face of uncertainty (OFU) principle [10]. By introducing $\tilde{P}\in\mathcal{M}$ as the additional decision variable, optimization (7) not only increases the chances of feasibility but also results in better regret bound [10]. However, as a result of this, optimization (7) now becomes complex to solve as $\tilde{P}$ is also a decision variable. The optimal value of the transition probability $\tilde{P}$ that maximizes optimization (7)-(8) is called the optimistic transition probability and denoted by $\tilde{P}_{k}$.
@@
If at any episode the extended LP (11) is not feasible, we apply a safe baseline policy. Once at an episode $k$, LP (11) is feasible, we demonstrate that the policy ${\pi}_{k}$ obtained using (13) is a safe policy with an arbitrarily high probability.
@@
We now derive an expression of a safe baseline policy with the knowledge of just a safe action for each proxy state and an upper bound of the stopping time of the MDP. In [15], we presented a safe baseline policy with the assumption that a deterministically safe action is available for each proxy state. However, this assumption is not realistic as a deterministically safe action may not always exist. This motivates us to design a safe baseline policy with a more general safe action namely $h$-safe action.
Theorem 8.
Consider $p\in(0,1)$ and an $h$-safe action $a^{s}(x)$ for each state $x$ with $h\leq\frac{p}{T_{max}}$. Then the following policy $\pi^{S}$, is a safe baseline policy for each state $x\in H\cup U$:
$\begin{split}\pi^{S}(a|x)=\begin{cases}\check{\pi}^{S}(a|x^{\prime}),&\text{if }x^{\prime}\in U^{\prime}\\ \text{random policy},&\text{if }x^{\prime}\in(H\setminus U^{\prime})\cup U,\end{cases}\end{split}$ (14)
where, for all $x^{\prime}\in U^{\prime}$,
$\begin{split}\check{\pi}^{S}(a|x)=\begin{cases}{q},&\text{if }a={a}^{s}(x^{\prime})\\ \frac{(1-q)}{|\mathcal{A}|-1},&\text{otherwise},\end{cases}\end{split}$
for any $q$ in the interval, $1\geq q\geq\frac{1-\frac{p}{T_{max}}}{1-h}$.
@@
Considering all the possible realizations, we obtain the equality $(a)$ as follows:
$\begin{split}&S^{P}_{\pi^{S}}(x_{0})=\mathds{E}_{\pi^{S}}^{x_{0}}[\sum_{t=0}^{\tau-1}\kappa_{t};P]\\ &\overset{}{\leq}\mathds{P}_{\pi^{S}}^{x_{0}}[\tau\leq T_{max};P]\mathds{E}_{\pi^{S}}^{x_{0}}[\sum_{t=0}^{T_{max}-1}\kappa_{t};P]\\ &+\mathds{P}_{\pi^{S}}^{x_{0}}[\tau>T_{max};P]\mathds{E}_{\pi^{S}}^{x_{0}}[\sum_{t=T_{max}}^{\infty}\kappa_{t};P]\end{split}$ (15)
@@
Furthermore,
$\mathds{E}_{\pi^{s}}^{x}\left[\kappa_{t};P\right]\leq\underset{x\in H\cup U}{\text{max}}\sum_{a\in A}\pi^{S}(a|x)\kappa(x,a),\ \forall t\in[0,T_{max})$
@@
$\begin{split}S^{P}_{\pi^{S}}(x_{0})&\overset{}{\leq}T_{max}\ \underset{t\in[0,T_{max})}{\text{max}}\ \mathds{E}_{\pi^{S}}^{x_{0}}\left[\kappa_{t};P\right]\\ &\overset{}{\leq}T_{max}\underset{{x}\in H\cup U}{\text{max}}\sum_{a\in A}\pi^{S}(a|x)\kappa(x,a)\\ &=T_{max}\underset{{x}\in H\cup U}{\text{max}}\sum_{y\in U}\Big[q\cdot P(x,a^{s}(x),y)\\ &+\sum_{a\in A\setminus\{a^{s}(x)\}}\frac{(1-q)}{|\mathcal{A}|-1}\cdot P(x,a,y)\Big]\\ &{=}T_{max}\Big[\underset{x\in H\cup U}{\text{max}}q\sum_{y\in U}P(x,a^{s}(x),y)\\ &+\underset{x\in H\cup U}{\text{max}}\sum_{a\in A\setminus\{a^{s}(x)\}}\frac{(1-q)}{|\mathcal{A}|-1}\sum_{y\in U}P(x,a,y)\Big]\end{split}$
@@
Equality $(e)$ follows from the definition of $h$-safe action. The following result demonstrates that the policy $\pi_{k}$ computed using the PSRL algorithm $1$ is safe with arbitrary high probability. The proof of the following theorem is similar to the proof of Theorem $2$ in [15].
Theorem 9.
@@
Theorem 11.
Consider $\delta\in(0,\frac{1}{2})$. With probability at least $(1-2\delta)$, PSRL has the following regret bound
$R(K)\overset{<}{\sim}\tilde{\mathcal{O}}\big(\frac{1}{p-p^{s}}|\mathcal{X}|\sqrt{|\mathcal{A}|T^{3}_{max}K}\big),$
where $p$ is the safety threshold, $p^{s}=S^{P}_{{\pi}^{s}}({x}_{0})$, $|\mathcal{X}|$ be the number of states, $|\mathcal{A}|$ be the number of actions, $T_{max}$ is the upper bound of the stopping time, and $K$ is the number of episodes.
@@ 4 Entropy regularized $p$-Safe RL (ER-PSRL) Algorithm
In this section, we present our main algorithm, entropy-regularized $p$-safe RL, which augments the underlying optimization problem with an entropy regularizer.
@@
Algorithm 1 suffers from two serious issues:
@@
Further, the entropy regularizer induces strong convexity in the optimization problem, ensuring that the optimal occupation measures depend continuously (Lipschitz continuously) on perturbations in both the empirical transition model and the cost coefficients. As a result, successive solutions vary smoothly across episodes, yielding stable policy updates and significantly reduced variability in the incurred cost and per-episode regret.
The entropy-regularized safe RL algorithm works almost exactly the same. In each episode $k\in\{1,2,...,K\}$, instead of computing the policy from Extended LP (11)-(12), we solve the following convex optimization.
@@
We define the entropy function $H$ with the following insight. From the definition of the occupation measure $\gamma(x,a)$ and $\beta(x,a)$ it is apparent that $\tilde{\xi}(x,a)\leq T_{max}$. We consider $\tilde{\xi}(x,a)/T_{max}$ instead of $\tilde{\xi}(x,a)$ as this will make $H\geq 0$ for all possible solutions $\tilde{g}(x,a)$ and $\tilde{h}(x,a)$. By introducing the parameter $\eta>0$ in the entropy function $H$ helps us in deriving the desired regret bound later.
@@
2:Initialize: Select $N_{1}(x,a)=0$, $N_{1}(x,a,y)=0$ for all $(x,a,y)\in H\times\mathcal{A}\times\mathcal{X}$, set $\alpha_{k}$ to some initial positive value.
@@
Theorem 12.
Consider $\delta\in(0,\frac{1}{2})$. With probability at least $(1-2\delta)$, the policy ${\pi}_{k}$ designed using the ER-PSRL algorithm is safe, i.e., the safety function $S^{P}_{{\pi}_{k}}({x}_{0})$ for the initial state $x_{0}$ satisfies the following with probability at least $(1-2\delta)$:
$S^{P}_{{\pi}_{k}}({x}_{0})\leq p.$
@@
Theorem 13.
Consider $\delta\in(0,\frac{1}{2})$. With probability at least $(1-2\delta)$, the ER-PSRL algorithm has the following regret bound.
$R(K)\overset{<}{\sim}\tilde{\mathcal{O}}\big(|\mathcal{X}|\sqrt{|\mathcal{A}|T^{3}_{max}K}\big),$
where $p$ is the safety threshold, $p^{s}=S^{P}_{{\pi}^{s}}({x}_{0})$, $|\mathcal{X}|$ be the number of states, $|\mathcal{A}|$ be the number of actions, $T_{max}$ is the upper bound of the stopping time, and $K$ is the number of episodes.
@@ 5 Simulation results
We consider an MDP with the following parameters. There are $5$ states given by $\mathcal{X}=\{1,2,3,4,5\}$ and two actions given by $\mathcal{A}=\{1,2\}$. The taboo set is $H=\{1,2,3\}$, unsafe state is $U=4$ and the goal state is $E=5$. The unsafe state and goal state are terminal. Clearly, the proxy set is $\{2,3\}$. State transition probabilities are as follows: $P(1,1,2)=0.9$, $P(1,1,3)=0.1$, $P(1,2,2)=0.1$, $P(1,2,3)=0.9$, $P(2,1,4)=0.8$, $P(2,1,5)=0.2$, $P(2,2,3)=0.2$, $P(2,2,5)=0.8$, $P(3,1,4)=0.8$, $P(3,1,5)=0.2$, $P(3,2,5)=1$. The remaining transition probabilities are 0. The safety parameter is $p=0.5$ and the confidence parameter is considered to be $\delta=0.01$. The cost associated with each state-action pair is as follows: c(1,1)=0.1c(1,1)=-0.1, c(1,2)=0.1c(1,2)=-0.1, c(2,1)=0.2c(2,1)=-0.2, c(2,2)=0.1c(2,2)=-0.1, c(3,1)=0.4c(3,1)=-0.4 and c(3,2)=0.1c(3,2)=-0.1. The optimal policy is given by $\pi^{*}(1|1)=0.46$, $\pi^{*}(2|1)=0.54$, $\pi^{*}(1|2)=0$, $\pi^{*}(2|2)=1.0$, $\pi^{*}(1|3)=1.0$ and $\pi^{*}(2|3)=0$ and the optimal value is $V_{\pi^{*}}^{P}=-0.396875$. We also assume that a deterministically safe action is available for each proxy state $2$ and $3$, which is Action $2$. Using Theorem 8, we design a safe baseline policy which is given by: $\pi(1|1)=\pi(2|1)=0.5$, $\pi^{s}(1|2)=\pi^{s}(1|3)=0.9$ and $\pi^{s}(2|2)=\pi^{s}(2|3)=0.9$. For state $1$, a uniformly random policy is used. The upper bound on the stopping time is considered to be $T_{max}=5$, which is clearly an upper bound of $\tau_{E}$ almost surely. We compute the safety function for the safe-baseline policy and is given by $p^{s}=S^{P}_{\pi^{s}}(1)=0.116$. The confidence parameter is considered to be $\delta=0.01$. In Figure 1 we show the per-episode objective regrets achieved by the PSRL algorithm and ER-PSRL algorithm. We clearly see that the per-episode objective regret suffered by PSRL algorithm varies quite drastically from one episode to another suggesting an unstable policy update. On the other hand, for the ER-PSRL algorithm, the per-episode objective regret does not deviate much, thus the ER-PSRL algorithm enjoys a stable and improves the policy consistently. Further, the ER-PSRL algorithm is also more efficient in learning better policy as we can see from Figure 2 that cumulative objective regret stay lower than the one with the PSRL algorithm. In Figure 3, we plot the per-episode constraint regret for both PSRL and ER-PSRL Algorithm. We see that for both algorithms, the per-episode constraint regret is below 0 suggesting that the safety constraint is always satisfied.
@@
We also compare the performance of the ER-PSRL algorithm under two scenarios: $(a)$ having the knowledge of the proxy set $U^{\prime}=\{2,3\}$ and $(b)$ not knowing the proxy set. If we know the proxy set, the safe baseline policy used until the feasibility of Optimization (16)-(12) is as given at the starting of this section. For the case when we do not know the proxy set $U^{\prime}$ then we consider the whole living set $H$ to be the proxy set. So the safe baseline policy is given by $\pi^{s}(1|1)=\pi^{s}(1|2)=\pi^{s}(1|3)=0.1$ and $\pi^{s}(2|1)=\pi^{s}(2|2)=\pi^{s}(2|3)=0.9$. In figure 4, we see that the knowledge of the proxy set results in lower cumulative objective regret, meaning that a better policy is learned. All of the simulation results are taken as the average of $5$ runs.

### v2

@@ 1 Introduction
To formalize this setup, we employ a CMDP formulation [3]. We assume that the target set is terminal, meaning that the process terminates whenever the MDP reaches a target state, while the unsafe set and the living set are non-terminal. In contrast, our earlier work [15] assumed that the unsafe set was also terminal in order to reduce computational complexity.
@@ 2 Background and Problem Formulation
For $t\leq\tau_{E}-1$, we define a sequence $\{c_{t}\}$ with $c_{t}:=c(X_{t},A_{t})$ and for all $t\geq\tau_{E}$, $c_{t}=0$. Similarly for $t\leq\tau-1$, we define another process $\{\kappa_{t}\}$ with $\kappa_{t}:=\kappa(X_{t},A_{t})$ and for all $t\geq\tau$, $\kappa_{t}:=0$.
Definition 1 ($p$-safety).
@@
If a safe baseline policy is enforced only on the proxy states while an arbitrary (possibly exploratory) policy is applied elsewhere, the overall safety of the system remains guaranteed [17]. Hence, the proxy set enables enhanced exploration by allowing less conservative actions without compromising safety. This work is, however, not restricted to the prior knowledge of the proxy set. In scenarios where the proxy set is unknown or cannot be identified apriori, we adopt a conservative assumption by treating all states as proxy states.
@@
The computed policy ${\pi}_{k}(a|x)$ is kept fixed till the end of the episode $k$. Optimization (7)-(8) is based on the optimization in the face of uncertainty (OFU) principle [10]. By introducing $\tilde{P}\in\mathcal{M}_{k}$ as the additional decision variable, optimization (7) not only increases the chances of feasibility but also results in better regret bound [10]. However, as a result of this, optimization (7) now becomes complex to solve as $\tilde{P}$ is also a decision variable. The optimal value of the transition probability $\tilde{P}$ that maximizes optimization (7)-(8) is called the optimistic transition probability and denoted by $\tilde{P}_{k}$.
@@
If at any episode the extended LP (11) is not feasible, we apply a safe baseline policy. Once at an episode $k$, LP (11) is feasible, we demonstrate that the policy ${\pi}_{k}$ obtained using (13) is a safe policy with an arbitrarily high probability. The pseudo-code for the PSRL algorithm is shown in Algorithm 1.
@@
We now derive a safe baseline policy using knowledge of a probabilistically safe action for each proxy state, together with an upper bound on the stopping time of the MDP. In our earlier work [15], the construction of a safe baseline policy relied on the assumption that a deterministically safe action is available at every proxy state. In practice, however, such deterministically safe actions may not exist. This observation motivates the development of a safe baseline policy that accommodates a more general notion of safety, wherein safe actions are characterized in a probabilistic sense.
Theorem 8 (Safe baseline policy).
Let $p\in(0,1)$ and suppose that for each state $x$ there exists a safe action $a^{s}(x)$ such that $h\leq\frac{p}{T_{\max}}$. Then, the policy $\pi^{S}$ defined below constitutes a safe baseline policy for all states $x\in H\cup U$:
$\pi^{S}(a\mid x)=\begin{cases}\check{\pi}^{S}(a\mid x^{\prime}),&\text{if }x^{\prime}\in U^{\prime},\\[5.69054pt] \text{uniform random policy},&\text{if }x^{\prime}\in(H\setminus U^{\prime})\cup U,\end{cases}$
where, for all $x^{\prime}\in U^{\prime}$, the policy $\check{\pi}^{S}$ is given by
$\check{\pi}^{S}(a\mid x^{\prime})=\begin{cases}q,&\text{if }a=a^{s}(x^{\prime}),\\[2.84526pt] \dfrac{1-q}{|\mathcal{A}|-1},&\text{otherwise},\end{cases}$
with the mixing parameter $q$ satisfying $1\geq q\geq\frac{1-\frac{p}{T_{\max}}}{1-h}$.
@@
Considering all the possible realizations, we obtain equality $(a)$ as follows:
$\begin{split}&S^{P}_{\pi^{S}}(x_{0})=\mathds{E}_{\pi^{S}}^{x_{0}}[\sum_{t=0}^{\tau-1}\kappa_{t};P]\\ &\overset{(a)}{\leq}\mathds{P}_{\pi^{S}}^{x_{0}}[\tau\leq T_{max};P]\mathds{E}_{\pi^{S}}^{x_{0}}[\sum_{t=0}^{T_{max}-1}\kappa_{t};P]\\ &+\mathds{P}_{\pi^{S}}^{x_{0}}[\tau>T_{max};P]\mathds{E}_{\pi^{S}}^{x_{0}}[\sum_{t=T_{max}}^{\infty}\kappa_{t};P]\end{split}$ (14)
@@
Furthermore, for all $x\in H$,
$\mathds{E}_{\pi^{s}}^{x}\left[\kappa_{t};P\right]\leq\underset{x\in H}{\text{max}}\sum_{a\in A}\pi^{S}(a|x)\kappa(x,a),\ \forall t\in[0,T_{max}-1]$
@@
$\begin{split}S^{P}_{\pi^{S}}(x_{0})&\overset{}{\leq}T_{max}\ \underset{t\in[0,T_{max}-1]}{\text{max}}\ \mathds{E}_{\pi^{S}}^{x_{0}}\left[\kappa_{t};P\right]\\ &\overset{}{\leq}T_{max}\ \underset{{x}\in H}{\text{max}}\sum_{a\in A}\pi^{S}(a|x)\kappa(x,a)\\ &=T_{max}\ \underset{{x}\in H}{\text{max}}\sum_{y\in U}\Big[q\cdot P(x,a^{s}(x),y)\\ &+\sum_{a\in A\setminus\{a^{s}(x)\}}\frac{(1-q)}{|\mathcal{A}|-1}\cdot P(x,a,y)\Big]\\ &{=}T_{max}\ \Big[\underset{x\in H}{\text{max}}\ q\ \sum_{y\in U}P(x,a^{s}(x),y)\\ &+\underset{x\in H}{\text{max}}\sum_{a\in A\setminus\{a^{s}(x)\}}\frac{(1-q)}{|\mathcal{A}|-1}\sum_{y\in U}P(x,a,y)\Big].\end{split}$
@@
Equality $(e)$ follows from the definition of the safe action. The following result shows that the policy $\pi_{k}$ computed by the PSRL algorithm is safe with arbitrarily high probability. The proof follows similar arguments to those used in the proof of Theorem $2$ in [15].
Theorem 9 (PSRL algorithm is $p$-safe).
@@
Theorem 11 (Regret bound of the PSRL algorithm).
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-2\delta)$, the cumulative objective regret of the PSRL algorithm after $K$ episodes satisfies
$R(K)\;\tilde{\mathcal{O}}\!\left(\frac{1}{p-p^{s}}\,|\mathcal{X}|\sqrt{|\mathcal{A}|\,T_{\max}^{3}\,K}\right),$
where, $p$ denotes the safety threshold, $p^{s}:=S^{P}_{\pi^{s}}(x_{0})$ is the safety level of the safe baseline policy $\pi^{s}$, $|\mathcal{X}|$ and $|\mathcal{A}|$ denote the numbers of states and actions, respectively, $T_{\max}$ is an upper bound on the stopping time, and $K$ is the number of episodes.
@@ 4 Entropy regularized $p$-Safe RL (ER-PSRL) Algorithm
In this section, we present our main algorithm, entropy-regularized $p$-safe reinforcement learning (ER–PSRL), which enhances the optimization problem underlying the PSRL algorithm by incorporating an entropy regularization term.
@@
The PSRL algorithm suffers from two serious issues:
@@
Further, the entropy regularizer induces strong convexity in the optimization problem, ensuring that the optimal occupation measures depend continuously on perturbations in both the empirical transition model $\hat{P}_{k}$. As a result, successive solutions vary smoothly across episodes, yielding stable policy updates and significantly reduced variability in the per-episode regret.
The ER-PSRL algorithm works almost exactly the same way as the PSRL algorithm. In each episode $k\in\{1,2,...,K\}$, instead of computing the policy from Extended LP (11)-(12), we solve the following convex optimization.
@@
We define the entropy function $H$ with the following insight. From the definition of the occupation measure $\gamma(x,a)$ and $\beta(x,a)$ it is apparent that $\tilde{\xi}(x,a)\leq T_{max}$ for all $(x,a)\in\mathcal{X}\times\mathcal{A}$. We consider $\tilde{\xi}(x,a)/T_{max}$ instead of $\tilde{\xi}(x,a)$ as this will make $H\geq 0$ for all possible solutions $\tilde{\xi}$. By introducing the parameter $\eta>0$ in the entropy function $H$ helps us in deriving the desired regret bound later.
@@
1:Input: Safety parameter $p\in(0,1)$, initial state $x_{0}$, confidence parameter $\delta\in(0,1)$, one safe action $a^{safe}(x^{\prime})$ for each proxy state, cost $c(x,a)$ for all $(x,a)\in\mathcal{X}\times\mathcal{A}$
@@
Theorem 12 (The ER–PSRL algorithm is $p$-safe).
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-2\delta)$, the policy $\pi_{k}$ computed by the ER–PSRL algorithm is safe, i.e., for the initial state $x_{0}$, the associated safety function satisfies
$S^{P}_{\pi_{k}}(x_{0})\leq p$
with probability at least $(1-2\delta)$.
@@
Theorem 13 (Regret bound of the ER-PSRL algorithm).
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-2\delta)$, the cumulative objective regret of the ER-PSRL algorithm after $K$ episodes satisfies
$R(K)\;\tilde{\mathcal{O}}\!\left(|\mathcal{X}|\sqrt{|\mathcal{A}|\,T_{\max}^{3}\,K}\right),$
where, $|\mathcal{X}|$ and $|\mathcal{A}|$ denote the numbers of states and actions, respectively, $T_{\max}$ is an upper bound on the stopping time, and $K$ is the number of episodes.
@@ 5 Simulation results
We consider a Markov decision process (MDP) defined as follows. The state space consists of five states, $\mathcal{X}=\{1,2,3,4,5\}$, and the action space contains two actions, $\mathcal{A}=\{1,2\}$. The living set is $H=\{1,2,3\}$, the unsafe state is $U=\{4\}$, and the goal state is $E=\{5\}$. Both the unsafe and the goal states are terminal. Clearly, the proxy state set is $\{2,3\}$. The nonzero state transition probabilities are given by P(1,1,2)=0.9,P(1,1,3)=0.1,P(1,2,2)=0.1,P(1,2,3)=0.9,P(2,1,4)=0.8,P(2,1,5)=0.2,P(2,2,3)=0.2,P(2,2,5)=0.8,P(3,1,4)=0.8,P(3,1,5)=0.2,P(3,2,5)=1.P(1,1,2)=0.9,\;P(1,1,3)=0.1,\;P(1,2,2)=0.1,\;P(1,2,3)=0.9,\;P(2,1,4)=0.8,\;P(2,1,5)=0.2,\;P(2,2,3)=0.2,\;P(2,2,5)=0.8,\;P(3,1,4)=0.8,\;P(3,1,5)=0.2,\;P(3,2,5)=1. All remaining transition probabilities are zero. The safety threshold is set to $p=0.5$, and the confidence parameter is chosen as $\delta=0.01$. The upper bound on the stopping time is chosen as $T_{\max}=5$, which almost surely upper bounds the hitting time $\tau_{E}$. The safety function corresponding to the safe baseline policy is computed as $p^{s}=S^{P}_{\pi^{s}}(1)=0.116$. The smoothing parameter is set to $\eta=0.1$. The cost associated with each state-action pair is given by c(1,1)=0.1c(1,1)=-0.1, c(1,2)=0.1c(1,2)=-0.1, c(2,1)=0.2c(2,1)=-0.2, c(2,2)=0.1c(2,2)=-0.1, c(3,1)=0.4c(3,1)=-0.4, and c(3,2)=0.1c(3,2)=-0.1. The optimal policy is $\pi^{*}(1|1)=0.46$, $\pi^{*}(2|1)=0.54$, $\pi^{*}(1|2)=0$, $\pi^{*}(2|2)=1$, $\pi^{*}(1|3)=1$, and $\pi^{*}(2|3)=0$, yielding an optimal value of $V^{P}_{\pi^{*}}=-0.396875$. We assume that a safe action is available at each proxy state $2$ and $3$, namely action $2$. Using Theorem 8, we construct a safe baseline policy $\pi^{s}$ given by $\pi^{s}(1|1)=\pi^{s}(2|1)=0.5$, $\pi^{s}(1|2)=\pi^{s}(1|3)=0.1$, and $\pi^{s}(2|2)=\pi^{s}(2|3)=0.9$. For state $1$, a uniformly random policy is employed.
Figure 1 illustrates the per-episode objective regret incurred by the PSRL and ER-PSRL algorithms. It can be observed that the per-episode objective regret of the PSRL algorithm exhibits significant variability across episodes, indicating unstable policy updates. In contrast, the per-episode objective regret of the ER-PSRL algorithm remains relatively stable, reflecting more consistent policy improvement.
Furthermore, the ER-PSRL algorithm demonstrates improved learning efficiency. As shown in Figure 2, its cumulative objective regret remains consistently lower than that of the PSRL algorithm. Finally, Figure 3 presents the per-episode constraint regret for both algorithms. For both PSRL and ER-PSRL, the per-episode constraint regret remains below zero, indicating that the safety constraint is satisfied at all times.
@@
We further compare the performance of the ER-PSRL algorithm under two scenarios: (i) when the proxy set $U^{\prime}=\{2,3\}$ is known, and (ii) when the proxy set is unknown. When the proxy set is known, the safe baseline policy employed until the optimization problem (15)–(12) becomes feasible is the one described at the beginning of this section. In contrast, when the proxy set $U^{\prime}$ is unknown, we conservatively treat the entire living set $H$ as the proxy set. In this case, the safe baseline policy is chosen as $\pi^{s}(1\mid 1)=\pi^{s}(1\mid 2)=\pi^{s}(1\mid 3)=0.1$ and $\pi^{s}(2\mid 1)=\pi^{s}(2\mid 2)=\pi^{s}(2\mid 3)=0.9$.
Figure 4 shows that having prior knowledge of the proxy set leads to a lower cumulative objective regret, indicating that a better policy is learned. All simulation results correspond to averages taken over $5$ independent runs.